OpenStudy (anonymous):

How many grams of HCl are produced when 15.2 mol H2O react?

OpenStudy (anonymous):

perhaps write the balanced equation to sus the stoichiometry

OpenStudy (anonymous):

HCl(g) + H2O(l) --> H3O+Cl-(aq) ( we see the molar ration governing HCL and H2O are 1:1 so number of moles of h20=number of moles of hcl

OpenStudy (anonymous):

n(hcl)=m/M m(hcl)=n*M m(hcl)=15.2*Molar mass find the molar mass and then multiply it with the number of moles( 15.2Moles) to find the mass of hcl in grams